The equation of the line that goes through the point ( 3 ,5 ) and is parallel to the line 2 x + 4 y = 2 can be written in the form y = mx+b where m is and where b is?
Answer #1
The slope of the line is -1/2:

2x + 4y = 2

4y = -2x + 2

y = -1/2 x + 1/2
answered by: ayinde
Answer #2
where m is the slop and b is y.so the equation in given with 2x+4y=2 and the points that is given is (3,5) so what you have to do is to put in this form y=mx+b so 3 is x value and 5 is y value.we have the equation for that witch is y-y1=m(x-x1)
y-5=2(x-3)
y-5=2x-6
y+5=2x+5
when you add 5 to both sides it will give you y=2x-1
answered by: saskshi